OpenStudy (anonymous):

PLEASE PLEASE PLEASE HELP :) Let f(x)=x2+6 . Find f(x+h)−f(x) All divided by (h) if h≠0 . Simplify your answer.

OpenStudy (anonymous):

f(x)=x2+6 f(x+h)=(x+h)2+6 :Simply replace x with x+h

OpenStudy (anonymous):

now you pretty much have everything you need to solve it

OpenStudy (anonymous):

can you help me solve it completely please, im really confused on how to solve it fully and then simplify it...

OpenStudy (mertsj):

\[\frac{(x+h)^2+6-(x^2+6)}{h}\]

OpenStudy (mertsj):

\[\frac{x^2+2xh+h^2+6-x^2-6}{h}\]

OpenStudy (mertsj):

\[\frac{2xh+h^2}{h}=2x+h\]

OpenStudy (anonymous):

how do you finish that simplification?

OpenStudy (mertsj):

it's done unless you are supposed to take the limit as h approaches 0 in which case it is 2x. Are you supposed to find the limit?

OpenStudy (mertsj):

And if you are, why didn't you post it?

OpenStudy (anonymous):

all it said was to find it if h is not equal to 0

OpenStudy (anonymous):

sorry the cross didnt go through the h :(

OpenStudy (mertsj):

Then the answer is 2x+h

OpenStudy (anonymous):

ah okay, well thank you!

OpenStudy (mikey):

@Mertsj.....why are you squaring x and (x+h)? The original function is 2x+6.

OpenStudy (mertsj):

because you wrote x2 and I thought you meant x^2. If you meant 2x, why didn't you write 2x instead of x2?

OpenStudy (mikey):

Ahh....that makes sense. :) I read it as 2x...but maybe he did leave out the ^.

OpenStudy (mertsj):

Who is "he"?

OpenStudy (mikey):

The poster of the problem.

OpenStudy (anonymous):

im a girl.